The campus suffered some lightning damage last night.
Beginning last night around 9:45 we started seeing problems.
Boliou Hall went off-line around 9:45PM. The cause was a failed UPS, which appears to have tripped out as a result of an over-voltage situation. All Boliou service was restored before 8AM this morning.
Goodsell Observatory and Old Music Hall suffered outages around 11PM. Upon investigation this morning, UPS devices in both buildings were operational, in each case one of the two Ethernet switches in the building were functioning normally.
Goodsell Observatory regained partial service around 8:15AM, with the full service restored by 9:15AM. One Ethernet switch was damaged. This was the switch connected to the outdoor telescope pedestals, providing a possible path for a lightening-induced surge.
Old Music Hall had one Ethenet switch damaged as well. All wireless service to Old Music Hall was not interrupted throughout the evening, but wired ports were down from 11PM until 8:15AM. I am currently by-passing the damaged switch module, but Old Music Hall will require a switch replacement in the near future. I will stage the cut over in such a way as to minimize downtime, but there could be user-visible disruptions of 30 seconds or so when the replacement equipment is cut in.
Farm House also suffered damage. Also, there may be related damage in Benton House and the ground floor of James Hall.
WiFi service has been restored to ground James and Benton House. Farm House has all services restored.
All users should be back to pre-storm service levels in all buildings. Pending remedial actions in Old Music Hall and James Hall may result in short localized outages.